OpenStudy (anonymous):

Assuming x > 0, what is the simplified fraction form for

OpenStudy (anonymous):

\[\frac{ 15x \sqrt{3} }{ \sqrt{75x^2}}\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Nurali @jhonyy9 @whpalmer4

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

can u help me @Nurali

OpenStudy (nurali):

\[\frac{ 15x \sqrt{3} }{ \sqrt{75x^2} }\]

OpenStudy (nurali):

\[\frac{ 15x \sqrt{3} }{ \sqrt{25*x^2*3} }\]

OpenStudy (nurali):

\[\frac{ 15x \sqrt{3} }{ 5x \sqrt{3}}\]

OpenStudy (nurali):

\[3\]

OpenStudy (anonymous):

ok thats it

OpenStudy (anonymous):

i appreciate it

OpenStudy (anonymous):

\[\frac{ 15x \sqrt{3} }{ 5x \sqrt{3} }\]

OpenStudy (anonymous):

is the last fraction form

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

No 3 is the last fraction

jhonyy9 (jhonyy9):

pardon but i think that like fraction form the last step will be 3/1